OpenStudy (anonymous):

1. Cameron had $500 in savings on January 1. Quinn had $800 in savings on January 1. Cameron deposits $20 per week into his savings account. Quinn withdraws $15 per week from his savings account. a) Write two equations: one for the amount of money in Cameron’s account x weeks after January 1st, and one for the amount of money in Quinn’s account x weeks after January 1st.

OpenStudy (anonymous):

hellpppppppppp

OpenStudy (anonymous):

second part to the question is 2. Use your answer from #1 to determine the number of weeks until Cameron will have more money in his savings account than Quinn. Show your work or provide an explanation for your answer.

OpenStudy (anonymous):

Cameron's amount starts at $500 and increases $20 per week. At week 0, he has 500. At week 1, he has 500 + 20(1 week). At week 2, he has 500 + 20(2weeks) = 500 +40.

OpenStudy (anonymous):

Does that give you any ideas? if C is Cameron's amount, it is going to be $500 (because of the initial balance) PLUS $20 per week. C = 500 + 20x when x is the number of weeks

OpenStudy (anonymous):

Quinn's is similar... he has $800 to start, but his DECREASES $15 for each week. Can you get the equation for Quinn?

OpenStudy (anonymous):

Q=800-15

OpenStudy (anonymous):

?

OpenStudy (anonymous):

close... that would be for the first week, right? Starts at 800, and after 1 week, it is 800 - 15 What about after 2 weeks?

OpenStudy (anonymous):

1600-15?

OpenStudy (anonymous):

i suck at math !

OpenStudy (anonymous):

well, maybe this will help a little. Are you saying that the 800 that Quinn started with somehow turns into $1600 after the second week? He is taking $15 out of his account each week. In week 1, he took out 15 from his original 800, so he has 800-15 = 785 left. In week 2, he took another 15 out from the 785 leftover, so it's 785-15 But notice... another way to talk about what happens in week 2 is that he started with 800 and took out 15 for two weeks in a row... 800-15-15. Each week, his balance gets smaller by 15... In week 3, he has 800 - 15 - 15 - 15 In week 4, he has 800 - 15 - 15 -15 - 15 and so on...

OpenStudy (anonymous):

it would be 785

OpenStudy (anonymous):

not 1600 i dont know why i put that i wanna take away 15 from 800

OpenStudy (anonymous):

yes, in the first week, it would be 785 What you need to do though is to find a "rule" that says what it will be every week. For the first part for Cameron, his rule was that he started with 500 and increased by 20 each week, or C = 500 + 20x . By putting in some number of "x" weeks, you can get Cameron's balance. For Quinn, he also has a rule, but it is a different rule: start at 800 and decrease by 15 dollars for each week. Q = 800 - 15x so for x = 1 week, Q = 800 - 15(1) = 800-15 = 785 and for week 10, Q = 800-15(10) = 800-150 = 650 and so on

OpenStudy (anonymous):

Does that make some sense? It might be a little confusing, but when they ask for the "equation of the amount of money in Cameron's account", they really just mean, "Please write down the rule that tells us how much he has for any number of weeks in the future"